the force between two charges is 125 N. One has a positive charge of 2.0 X 10-3 C, while the other has negative charge of 3.6 X 10-4 C. How far apart are the two charges?


1
Expert's answer
2020-07-21T13:16:13-0400

According to coulombs law

"f=\frac{k\times Q_{1} \times Q_{2}}{r^{2}}" where "k=9.0\times 10^{9}Nm^{2}/c^{2}"


therefore "r^{2}= \frac{k\times Q_{1}\times Q_{2}}{f}"

substituting the given values in the equation


"r^{2}=\frac {(9.0\times 10^{9})\times(2.0\times 10^{-3})\times(3.6\times10^{-4})}{125}"

"r^{2}=51.84"


"r=\sqrt{51.84}"

.

"r=7.2m"